OpenStudy (anonymous):

i am Stu6id so pls help thanks love yah Solve for x: 3 over 4 x + 5 over 8 = 4x

OpenStudy (chantysquirrel1129**):

Need halp? :3

zepdrix (zepdrix):

\[\large\rm \frac{3}{4x}+\frac{5}{8}=4x\]

OpenStudy (chantysquirrel1129**):

or is it: \[\frac{ 3 }{ 4 }x+\frac{ 5 }{ 8 }=4x\] :o

OpenStudy (anonymous):

ok so wat exactly do I do

OpenStudy (chantysquirrel1129**):

Which equation is your algebra problem, because it does make a difference :)

OpenStudy (anonymous):

i already got it thats ppl

OpenStudy (anonymous):

oh and for future people its 1/6

OpenStudy (chantysquirrel1129**):

Lol oh cx xD